# Quantum metrics from length functions on quantum groups

### v1

@@ 3. Dirac operators from length functions
This section is devoted to a study of the spectral triples arising from a length function on a compact quantum group. Length functions on classical discrete groups have played a prominent role in non-commutative geometry since Connes’ seminal paper [Con89], and have also been studied in detail from the quantum metric point of view in [OR05, CR17, Rie02]. Length functions have also appeared in the quantum group literature in connection with the rapid decay property [Ver07, BVZ15]. We recall the definition here:
@@
Let ${\mathbb{G}}$ be a compact quantum group. A length function on ${\mathbb{G}}$ is a function $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$ such that
@@
From a proper length function, one obtains an unbounded operator $\mathcal{D}_{\ell}\colon\Lambda(\operatorname{Pol}({\mathbb{G}}))\to L^{2}({\mathbb{G}})$ by setting
$\mathcal{D}_{\ell}(\Lambda(u_{ij}^{\alpha})):=\ell(u^{\alpha})\cdot\Lambda(u_{ij}^{\alpha}).$ (3.1)
As the set $\{\tfrac{1}{\|\Lambda(u_{ij}^{\alpha})\|}\Lambda(u^{\alpha}_{ij})\}$ forms an orthonormal basis for $L^{2}({\mathbb{G}})$, this defines an essentially self-adjoint operator and we denote its self-adjoint closure by $D_{\ell}$. In analogy with the case of classical discrete groups, one obtains a non-commutative geometry in this way:
Lemma 3.2 ([BVZ15, Lemma 5.4]).
@@
As already mentioned in the introduction, the main purpose of the present paper is to investigate the quantum metric structure associated with the commutator seminorm arising from this construction. This idea was already investigated in [BVZ15], where the authors prove that in the presence of the rapid decay property, one does indeed obtain a compact quantum metric structure by taking iterated commutators with $D_{\ell}$, where the number of iterations depends on the degree of rapid decay. To the best of our knowledge, [BVZ15, Theorem 7.4] is the only positive result in this direction, but even for reasonably tame examples, such as $SU_{q}(2)$, their result seems to require taking at least two commutators before yielding a Lip-norm. In particular, there are no known examples of length functions on (genuine) quantum groups for which the basic commutator seminorm gives a compact quantum metric structure. Iterated commutator seminorms have the drawback that they will rarely satisfy the Leibniz rule, thus straying from the original noncommutative geometric motivation behind the problem. The aim in the present paper is to provide the first examples for which a single commutator suffices; see Corollary C. To this end, the fusion algebra will play a central role and in the following section we will show how one may also base a spectral triple on this algebra.
@@ 3.1. A spectral triple for the fusion algebra
Consider again a compact quantum group ${\mathbb{G}}$ and its fusion algebra $F({\mathbb{G}})=\textup{span}_{{\mathbb{C}}}\{\alpha\mid\alpha\in\operatorname{Irred}({\mathbb{G}})\}$. We endow $F({\mathbb{G}})$ with the inner product for which the elements in $\operatorname{Irred}({\mathbb{G}})$ form an orthonormal basis, and denote the completion by $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$, and by $c_{c}(\operatorname{Irred}({\mathbb{G}}))$ the dense subspace spanned by the orthonormal basis; i.e. the image of $F({\mathbb{G}})$ in $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$. For each $x\in F({\mathbb{G}})$, we therefore obtain a linear map ${\pi}_{0}(x)\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\to c_{c}(\operatorname{Irred}({\mathbb{G}}))$ which multiplies from the left with $x$, and we now show that this map extends boundedly to the Hilbert space $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$. Recall that $\lambda$ denotes the GNS representation of $C({\mathbb{G}})$ on $L^{2}({\mathbb{G}})$.
@@
The character map $\chi\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\to\Lambda(\operatorname{Pol}_{z}({\mathbb{G}}))$, $\alpha\mapsto\sum_{i=1}^{d_{\alpha}}\Lambda(u_{ii}^{\alpha})$, extends to a unitary $\tilde{\chi}\colon\ell^{2}(\operatorname{Irred}({\mathbb{G}}))\to L^{2}_{z}({\mathbb{G}}):=\overline{\Lambda(\operatorname{Pol}_{z}(({\mathbb{G}}))}\subset L^{2}({\mathbb{G}})$. For $x\in F({\mathbb{G}})$, it holds that
@@
as operators on $c_{c}(\operatorname{Irred}({\mathbb{G}}))$, and ${\pi}_{0}(x)$ therefore extends to a bounded operator ${\pi}(x)$ on $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$ which equals $\tilde{\chi}^{*}\pi(\chi(x))\tilde{\chi}$.
@@
By definition, the image of $\chi$ is a dense subspace in $L^{2}_{z}({\mathbb{G}})$. Moreover, by the Schur orthogonality relations [Tim08, Proposition 5.3.8 (iii)], we obtain that that
$\left\langle\chi({\alpha}),\chi(\beta)\right\rangle=\sum_{i=1}^{d_{\alpha}}\sum_{j=1}^{d_{\beta}}h(u_{ii}^{\alpha}(u_{jj}^{\beta})^{*})=\delta_{\alpha,\beta}\sum_{i=1}^{d_{\alpha}}h(u_{ii}^{\alpha}(u_{ii}^{\alpha})^{*})=\delta_{\alpha,\beta}.$
@@
The Haar state, $h$, pulls back to a positive functional $\tau\colon F({\mathbb{G}})\to{\mathbb{C}}$ via the character map $\chi$, which on the basis $\operatorname{Irred}({\mathbb{G}})$ is given by $\tau(\alpha)=\delta_{\alpha,e}$. By Frobenius reciprocity, the map $\tau$ is a trace and hence so is the restriction of $h$ to $\operatorname{Pol}_{z}({\mathbb{G}})$. Also note that the representation ${\pi}$ constructed above agrees with the GNS representation arising from $\tau$, and the latter therefore, in particular, gives rise to bounded operators; see also [HI98] for a more general approach to this.
We now furthermore assume that ${\mathbb{G}}$ is endowed with a proper length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$. Setting $A_{n}:=\textup{span}\{\alpha\mid\ell(\alpha)\leqslant n\}$ provides a filtration of $F({\mathbb{G}})$, in the sense that $A_{0}={\mathbb{C}}\cdot e$, $A_{n}^{*}=A_{n}$ and $A_{n}\cdot A_{m}\subseteq A_{n+m}$, which follows directly from the defining properties of $\ell$. These comments, together with Remark 3.4, show that the fusion algebra $F({\mathbb{G}})$ falls within the class of filtered *-algebras studied by Ozawa and Rieffel in [OR05].
The length function also defines an unbounded operator $\tilde{\mathcal{D}}_{\ell}\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\to\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$ given by $\tilde{\mathcal{D}}_{\ell}(\alpha)=\ell(\alpha)\cdot\alpha$. This operator is essentially self-adjoint, and we denote its self-adjoint closure by $\tilde{{D}}_{\ell}$.
@@
The data $\big{(}F({\mathbb{G}}),\ell^{2}(\operatorname{Irred}({\mathbb{G}})),\tilde{D}_{\ell}\big{)}$ is a spectral triple.
@@
As noted above, our setup is compatible with that studied in [OR05] and the fact that $\tilde{D}_{\ell}$ has bounded commutators with elements from $F({\mathbb{G}})$ therefore follows from [OR05, Lemma 1.1]. That $\tilde{D}_{\ell}$ has compact resolvent is trivial if $\operatorname{Irred}({\mathbb{G}})$ is finite and follows from the properness of $\ell$ when $\operatorname{Irred}({\mathbb{G}})$ is infinite. ∎
@@
The spectral triples $\big{(}F({\mathbb{G}}),\ell^{2}(\operatorname{Irred}({\mathbb{G}})),\tilde{D}_{\ell}\big{)}$ and $\big{(}\operatorname{Pol}_{z}({\mathbb{G}}),L^{2}_{z}({\mathbb{G}}),D_{\ell}^{z}\big{)}$ are unitarily equivalent via the unitary $\tilde{\chi}$. In particular, if one is a spectral metric space, then so is the other.
@@
If $\big{(}F({\mathbb{G}}),\ell^{2}(\operatorname{Irred}({\mathbb{G}})),\tilde{D}_{\ell}\big{)}$ is a spectral metric space then the restriction of $L_{\ell}$ provides $C_{z}({\mathbb{G}})$ with the structure of a compact quantum metric space.
@@
By assumption, $\big{(}F({\mathbb{G}}),\ell^{2}(\operatorname{Irred}({\mathbb{G}})),\tilde{D}_{\ell}\big{)}$ is a spectral metric space and by Proposition 3.6 the same is therefore true for $(\operatorname{Pol}_{z}({\mathbb{G}}),L^{2}_{z}({\mathbb{G}}),D_{\ell}^{z})$. By Rieffel’s criterion (Theorem 2.6), this is equivalent to the Lipschitz unit ball $\{a\in\operatorname{Pol}_{z}({\mathbb{G}})\mid\|[D_{\ell}^{z},a]\|\leqslant 1\}$ having pre-compact image in $C_{z}({\mathbb{G}})/{\mathbb{C}}$. By the inequality established above,
@@ 4. The conjugation coaction
[…]
@@
The key tool will be the conjugation coaction studied in [AC17, Cra19], and we will therefore be following their conventions below. It is not hard to see that both $W$ and $\Sigma V\Sigma$ are unitary corepresentations. From (2.5), it follows that $W_{23}$ and $(\Sigma V\Sigma)_{13}$ commute and hence also $Z:=W\Sigma V\Sigma\in L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(L^{2}({\mathbb{G}}))$ becomes a unitary corepresentation. The associated coaction $\delta\colon{\mathbb{B}}(L^{2}({\mathbb{G}}))\to L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(L^{2}({\mathbb{G}}))$,
@@
is going to be the quantum analogue of the conjugation action on a classical group discussed above, and we will therefore refer to it as the conjuguation coaction. We denote by ${\mathbb{B}}(L^{2}({\mathbb{G}}))^{{\mathbb{G}}}$ the corresponding fixed point algebra
$\{a\in{\mathbb{B}}(L^{2}({\mathbb{G}}))\mid\delta(T)=1\otimes T\}.$
@@
By [Wan17, Lemma 6.3], there exists an $h$-preserving conditional expectation from $C({\mathbb{G}})$ onto $C_{z}({\mathbb{G}})$ exactly when ${\mathbb{G}}$ is of Kac type. In order to show that this conditional expectation is a contraction for seminorms of the form $L_{\ell}$, we will need a quantum analogue of the formula (4.1), which is obtained in the following proposition.
@@
The map $E\colon{\mathbb{B}}(L^{2}({\mathbb{G}}))\to{\mathbb{B}}(L^{2}({\mathbb{G}}))$ given by $E(a):=(h\otimes{\operatorname{id}})\delta(a)$ defines a normal conditional expectation onto the fixed point algebra ${\mathbb{B}}(L^{2}({\mathbb{G}}))^{{\mathbb{G}}}$, which preserves the vector state $h$ implemented by $\Lambda(1)$. If ${\mathbb{G}}$ is of Kac type, $E$ maps $L^{\infty}({\mathbb{G}})$ to $L^{\infty}_{z}({\mathbb{G}})$, $C({\mathbb{G}})$ to $C_{z}({\mathbb{G}})$ and $\operatorname{Pol}({\mathbb{G}})$ to $\operatorname{Pol}_{z}({\mathbb{G}})$.
@@
The map $E$ is unital and completely positive (henceforth abbreviated ucp) by construction. Moreover, $T\mapsto 1\otimes T$ is normal as a *-isomorphism of ${\mathbb{B}}(L^{2}({\mathbb{G}}))$ onto the von Neumann algebra $1\otimes{\mathbb{B}}(L^{2}({\mathbb{G}}))$ [Bla06, Section III, Corollary 2.2.12]. From this it follows that $\delta$ is normal and since also ${\operatorname{id}}\otimes h$ is normal, we conclude that $E$ is normal.
@@
$h(E(a))$ $=h(({\operatorname{id}}\otimes h)\delta(a))$
@@
It only remains to be shown that $E$ preserves the algebras associated with ${\mathbb{G}}$ in the Kac type situation. Arguing as in [Lem15, proof of Theorem 1.4], we first note that for $x\in C({\mathbb{G}})$ we have
@@
and $\Delta$ therefore extends isometrically to $\tilde{\Delta}\colon L^{2}({\mathbb{G}})\to L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$. For $a\in C({\mathbb{G}})$, we now claim that
@@
$\left\langle E(a)\Lambda(x),\Lambda(y)\right\rangle$ $=\left\langle(h\otimes 1)\delta(a)\Lambda(x),\Lambda(y)\right\rangle$
$=\left\langle\delta(a)(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle$
$=\left\langle\Sigma V^{*}\Sigma W^{*}(1\otimes a)W\Sigma V\Sigma(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle$
$=\left\langle\Sigma\Delta(a)\Sigma V(\Lambda(x)\otimes\Lambda(1)),V(\Lambda(y)\otimes\Lambda(1))\right\rangle$
$=\left\langle\Sigma\Delta(a)\Sigma(\Lambda\otimes\Lambda(\Delta(x))),\Lambda\otimes\Lambda(\Delta(y))\right\rangle$
$=\left\langle\Delta^{\operatorname{op}}(a)\circ\tilde{\Delta}(\Lambda(x)),\tilde{\Delta}(\Lambda(y))\right\rangle$
$=\left\langle\tilde{\Delta}^{*}\circ\Delta^{\operatorname{op}}(a)\circ\tilde{\Delta}(\Lambda(x)),\Lambda(y)\right\rangle.$
Using the formula (4.2), we will now show that when ${\mathbb{G}}$ is of Kac type and $a\in\operatorname{Pol}({\mathbb{G}})$ then $E(a)\in L^{\infty}({\mathbb{G}})$. From this it follows that $E$ preserves $L^{\infty}({\mathbb{G}})$, since $E$ is normal and $\operatorname{Pol}({\mathbb{G}})$ is ultraweakly dense in $L^{\infty}({\mathbb{G}})$, and hence that $E$ maps $L^{\infty}({\mathbb{G}})$ to $L^{\infty}_{z}({\mathbb{G}})$ by Lemma 4.1. Since $h$ is a trace, the map $\Lambda(x)\to\Lambda(x^{*})$ extends to an anti-unitary $J$ on $L^{2}({\mathbb{G}})$ with the property that $JL^{\infty}({\mathbb{G}})J=L^{\infty}({\mathbb{G}})^{\prime}$. By von Neumann’s bicommutant theorem and density of $\operatorname{Pol}({\mathbb{G}})$ in $L^{\infty}({\mathbb{G}})$, it therefore suffices to show that $E(a)$ commutes with $JbJ$ for all $b\in\operatorname{Pol}({\mathbb{G}})$. To see this, fix again $x,y\in\operatorname{Pol}({\mathbb{G}})$ and note that $JbJ\Lambda(x)=\Lambda(xb^{*})$. Using (4.2), we now compute as follows:
@@
$=\left\langle\Lambda\otimes\Lambda(a_{(2)}x_{(1)}\otimes a_{(1)}x_{(2)}),\Lambda\otimes\Lambda(y_{(1)}b_{(1)})\otimes(y_{(2)}b_{(2)})\right\rangle$

### v3

@@ 3. Dirac operators from length functions
This section is devoted to a study of the spectral triples arising from a length function on a compact quantum group. Length functions on classical discrete groups have played a prominent role in non-commutative geometry since Connes’ seminal paper [Con:CFH], and have also been studied in detail from the quantum metric point of view in [OzawaRieffel2005, ChristRieffel2017, Rie:GrpCstar]. Length functions have also appeared in the quantum group literature in connection with the rapid decay property [Vergnioux:rapid, BVZ:rapid-decay]. We recall the definition here:
@@
Let $\mathbb{G}$} be a compact quantum group. A length function on $\mathbb{G}$} is a function $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$ such that
@@
The standard example of a length function, and indeed the only one for which we will provide examples in Section 6, arises when $\operatorname{Irred}({\mathbb{G}})$ is finitely generated.  This, by definition, means that there exists a finite set $S\subset\operatorname{Irred}({\mathbb{G}})$ such that every $\alpha\in\operatorname{Irred}({\mathbb{G}})$ is equivalent to a sub-corepresentation in a tensor product of elements from $S$. Setting $\ell(\alpha)$ equal to the smallest integer $k$ such that $\alpha$ is equivalent to a sub-corepresentation of a tensor product of $k$ elements from $S$ provides a length function on $\mathbb{G}$}, commonly referred to as the word length function (with respect to $S$). In Section 6, our primary interest will be in the situation where $S$ can be chosen to consist of single element $u$ (this is for instance the case for quantum $SU(2)$), and in this situation $\ell(\alpha)$ is simply the smallest $k$ for which $\alpha$ appears in the decomposition of the tensor power $u^{\boxtimes k}$.  For this reason, and to formally conform with the setup in [OzawaRieffel2005], we will now restrict to proper length functions with values in ${\mathbb{N}}_{0}$.
From a proper length function, one obtains an unbounded operator $\mathcal{D}_{\ell}\colon\Lambda(\operatorname{Pol}({\mathbb{G}}))\rightarrow L^{2}({\mathbb{G}})$ by setting
$\mathcal{D}_{\ell}(\Lambda(u_{ij}^{\alpha})):=\ell(\alpha)\cdot\Lambda(u_{ij}^{\alpha}).$ (3.1)
As the set $\{\tfrac{1}{\|\Lambda(u_{ij}^{\alpha})\|_{2}}\Lambda(u^{\alpha}_{ij})\}$ forms an orthonormal basis for $L^{2}({\mathbb{G}})$, this defines an essentially self-adjoint operator and we denote its self-adjoint closure by $D_{\ell}$. In analogy with the case of classical discrete groups, one obtains a non-commutative geometry in this way. Note that the following result is stated only for word length functions in [BVZ:rapid-decay, Lemma 5.4], but their proof only depends on [OzawaRieffel2005, Lemma 1.1], so it also holds true in our setting.
Lemma 3.2 ([BVZ:rapid-decay, Lemma 5.4]).
@@
As already mentioned in the introduction, the main purpose of the present paper is to investigate the quantum metric structure associated with the commutator seminorm arising from this construction. This idea was already investigated in [BVZ:rapid-decay], where the authors prove that in the presence of the rapid decay property, one does indeed obtain a compact quantum metric structure by taking iterated commutators with $D_{\ell}$, where the number of iterations depends on the degree of rapid decay. To the best of our knowledge, [BVZ:rapid-decay, Theorem 7.4] is the only positive result in this direction, but even for reasonably tame examples, such as $SU_{q}(2)$, their result seems to require taking at least two commutators before yielding a Lip-norm. In particular, there are no known examples of length functions on (genuine) quantum groups for which the basic commutator seminorm gives a compact quantum metric structure. Iterated commutator seminorms have the drawback that they will rarely satisfy the Leibniz rule, thus straying from the original noncommutative geometric motivation behind the problem. The aim in the present paper is to provide the first examples for which a single commutator suffices; see Corollary C. To this end, the fusion algebra will play a central role and in the following section we will show how one may also base a spectral triple on this algebra.
@@ 3.1. A spectral triple for the fusion algebra
Consider again a compact quantum group $\mathbb{G}$} and its fusion algebra $F({\mathbb{G}})=\textup{span}_{{\mathbb{C}}}\{\alpha\mid\alpha\in\operatorname{Irred}({\mathbb{G}})\}$. We endow $F({\mathbb{G}})$ with the inner product for which the elements in $\operatorname{Irred}({\mathbb{G}})$ form an orthonormal basis, and denote the completion by $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$, and by $c_{c}(\operatorname{Irred}({\mathbb{G}}))$ the dense subspace spanned by the orthonormal basis; i.e. the image of $F({\mathbb{G}})$ in $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$. For each $x\in F({\mathbb{G}})$, we therefore obtain a linear map ${\pi}_{0}(x)\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\rightarrow c_{c}(\operatorname{Irred}({\mathbb{G}}))$ which multiplies from the left with $x$, and we now show that this map extends boundedly to the Hilbert space $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$. Recall that $\lambda$ denotes the GNS representation of $C({\mathbb{G}})$ on $L^{2}({\mathbb{G}})$.
@@
The character map $\chi\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\rightarrow\Lambda(\operatorname{Pol}_{z}({\mathbb{G}}))$, $\alpha\mapstochar\rightarrow\sum_{i=1}^{d_{\alpha}}\Lambda(u_{ii}^{\alpha})$, extends to a unitary $\tilde{\chi}\colon\ell^{2}(\operatorname{Irred}({\mathbb{G}}))\rightarrow L^{2}_{z}({\mathbb{G}}):=\overline{\Lambda(\operatorname{Pol}_{z}(({\mathbb{G}}))}\subset L^{2}({\mathbb{G}})$. For $x\in F({\mathbb{G}})$, it holds that
@@
as operators on $c_{c}(\operatorname{Irred}({\mathbb{G}}))$, and ${\pi}_{0}(x)$ therefore extends to a bounded operator ${\pi}(x)$ on $\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$ which equals $\tilde{\chi}^{*}\lambda(\chi(x))\tilde{\chi}$.
@@
By definition, the image of $\chi$ is a dense subspace in $L^{2}_{z}({\mathbb{G}})$. Moreover, by the Schur orthogonality relations [Timmermann-book, Proposition 5.3.8 (iii)], we obtain that
$\left\langle\chi({\alpha}),\chi(\beta)\right\rangle=\sum_{i=1}^{d_{\alpha}}\sum_{j=1}^{d_{\beta}}h((u_{ii}^{\alpha})^{*}u_{jj}^{\beta}))=\delta_{\alpha,\beta}\sum_{i=1}^{d_{\alpha}}h((u_{ii}^{\alpha})^{*}u_{ii}^{\alpha})=\delta_{\alpha,\beta}.$
@@
The Haar state, $h$, pulls back to a positive functional $\tau\colon F({\mathbb{G}})\rightarrow{\mathbb{C}}$ via the character map $\chi$, which on the basis $\operatorname{Irred}({\mathbb{G}})$ is given by $\tau(\alpha)=\delta_{\alpha,e}$. By Frobenius reciprocity, the map $\tau$ is a trace and hence so is the restriction of $h$ to $\operatorname{Pol}_{z}({\mathbb{G}})$. Also note that the representation $\pi$} constructed above agrees with the GNS representation arising from $\tau$, and the latter therefore, in particular, gives rise to bounded operators; see also [Hiai-Izumi] for a more general approach to this.
We now furthermore assume that $\mathbb{G}$} is endowed with a proper length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow{\mathbb{N}}_{0}$. Setting $A_{n}:=\textup{span}\{\alpha\mid\ell(\alpha)\leqslant n\}$ provides a filtration of $F({\mathbb{G}})$, in the sense that $A_{0}={\mathbb{C}}\cdot e$, $A_{n}^{*}=A_{n}$ and $A_{n}\cdot A_{m}\subseteq A_{n+m}$, which follows directly from the defining properties of $\ell$. These comments, together with Remark 3.4, show that the fusion algebra $F({\mathbb{G}})$ falls within the class of filtered *-algebras studied by Ozawa and Rieffel in [OzawaRieffel2005].
The length function also defines an unbounded operator $\tilde{\mathcal{D}}_{\ell}\colon c_{c}(\operatorname{Irred}({\mathbb{G}}))\rightarrow\ell^{2}(\operatorname{Irred}({\mathbb{G}}))$ given by $\tilde{\mathcal{D}}_{\ell}(\alpha)=\ell(\alpha)\cdot\alpha$. This operator is essentially self-adjoint, and we denote its self-adjoint closure by $\tilde{{D}}_{\ell}$.
@@
[…]
@@
As noted above, our setup is compatible with that studied in [OzawaRieffel2005] and the fact that $\tilde{D}_{\ell}$ has bounded commutators with elements from $F({\mathbb{G}})$ therefore follows from [OzawaRieffel2005, Lemma 1.1]. That $\tilde{D}_{\ell}$ has compact resolvent is trivial if $\operatorname{Irred}({\mathbb{G}})$ is finite and follows from the properness of $\ell$ when $\operatorname{Irred}({\mathbb{G}})$ is infinite. ∎
@@
[…]
@@
[…]
@@
[…]
@@ 4. The conjugation coaction
For a classical compact group $G$, the algebra $C_{z}(G)$ consists of the conjugation invariant functions, i.e. those $f\in C(G)$ for which $f(g^{-1}xg)=f(x)$ for all $g,x\in G$. The algebra $C_{z}(G)$ may therefore be viewed as the fixed point algebra for the conjugation action $G\overset{\alpha}{\curvearrowright}C(G)$, and one obtains a natural conditional expectation $E\colon C(G)\rightarrow C_{z}(G)$ by integrating the action against the Haar probability measure $\mu$ on $G$:
$E(f)=\int_{G}\alpha_{g}(f)\ d\mu(g)$ (4.1)
@@
The key tool will be the conjugation coaction studied in [Crann:CaracterDensity, Crann:InnerAmenability], and we will therefore be following their conventions below. It is not hard to see that both $W$ and $\Sigma V\Sigma$ are unitary corepresentations, where $W$ and $V$ are the fundamental unitaries from Section 2.2.1. From (2.5), it follows that $W_{23}$ and $(\Sigma V\Sigma)_{13}$ commute and hence also $Z:=W\Sigma V\Sigma\in L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(L^{2}({\mathbb{G}}))$ becomes a unitary corepresentation. The associated coaction $\delta\colon{\mathbb{B}}(L^{2}({\mathbb{G}}))\rightarrow L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(L^{2}({\mathbb{G}}))$,
@@
is going to be the quantum analogue of the conjugation action on a classical group discussed above, and we will therefore refer to it as the conjugation coaction. We denote by ${\mathbb{B}}(L^{2}({\mathbb{G}}))^{{\mathbb{G}}}$ the corresponding fixed point algebra
$\{T\in{\mathbb{B}}(L^{2}({\mathbb{G}}))\mid\delta(T)=1\otimes T\}.$
@@
By [Wang:lacunary, Lemma 6.3], there exists an $h$-preserving conditional expectation from $C({\mathbb{G}})$ onto $C_{z}({\mathbb{G}})$ exactly when $\mathbb{G}$} is of Kac type. In order to show that this conditional expectation is a contraction for seminorms of the form $L_{\ell}$, we will need a quantum analogue of the formula (4.1), which is obtained in the following proposition.
@@
The map $E\colon{\mathbb{B}}(L^{2}({\mathbb{G}}))\rightarrow{\mathbb{B}}(L^{2}({\mathbb{G}}))$ given by $E(a):=(h\otimes{\operatorname{id}})\delta(a)$ defines a normal conditional expectation onto the fixed point algebra ${\mathbb{B}}(L^{2}({\mathbb{G}}))^{{\mathbb{G}}}$, which preserves the vector state $h$ implemented by $\Lambda(1)$. If $\mathbb{G}$} is of Kac type, $E$ maps $L^{\infty}({\mathbb{G}})$ to $L^{\infty}_{z}({\mathbb{G}})$, $C({\mathbb{G}})$ to $C_{z}({\mathbb{G}})$ and $\operatorname{Pol}({\mathbb{G}})$ to $\operatorname{Pol}_{z}({\mathbb{G}})$.
@@
The map $E$ is unital and completely positive (henceforth abbreviated ucp) by construction. Moreover, $T\mapstochar\rightarrow 1\otimes T$ is normal as a *-isomorphism of ${\mathbb{B}}(L^{2}({\mathbb{G}}))$ onto the von Neumann algebra $1\otimes{\mathbb{B}}(L^{2}({\mathbb{G}}))$ [Blackadar:OperatorAlgebras, Section III, Corollary 2.2.12]. From this it follows that $\delta$ is normal and since also $h\otimes{\operatorname{id}}$ is normal, we conclude that $E$ is normal.
@@
$h(E(a))$ $=h((h\otimes{\operatorname{id}})\delta(a))$
@@
It only remains to be shown that $E$ preserves the algebras associated with $\mathbb{G}$} in the Kac type situation. Arguing as in [Lemeux:Haagerup, proof of Theorem 1.4], we first note that for $x\in C({\mathbb{G}})$ we have
@@
and $\Delta$ therefore extends isometrically to $\tilde{\Delta}\colon L^{2}({\mathbb{G}})\rightarrow L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$. For $a\in C({\mathbb{G}})$, we now claim that
@@
$\left\langle\Lambda(x),E(a)\Lambda(y)\right\rangle$ $=\left\langle\Lambda(x),(h\otimes 1)\delta(a)\Lambda(y)\right\rangle$
$=\left\langle\Lambda(1)\otimes\Lambda(x),\delta(a)(\Lambda(1)\otimes\Lambda(y))\right\rangle$
$=\left\langle\Lambda(1)\otimes\Lambda(x),\Sigma V^{*}\Sigma W^{*}(1\otimes a)W\Sigma V\Sigma(\Lambda(1)\otimes\Lambda(y))\right\rangle$
$=\left\langle V(\Lambda(x)\otimes\Lambda(1)),\Sigma\Delta(a)\Sigma V(\Lambda(y)\otimes\Lambda(1))\right\rangle$
$=\left\langle\Lambda\otimes\Lambda(\Delta(x)),\Sigma\Delta(a)\Sigma(\Lambda\otimes\Lambda(\Delta(y)))\right\rangle$
$=\left\langle\tilde{\Delta}(\Lambda(x)),\Delta^{\operatorname{op}}(a)\circ\tilde{\Delta}(\Lambda(y))\right\rangle$
$=\left\langle\Lambda(x),\tilde{\Delta}^{*}\circ\Delta^{\operatorname{op}}(a)\circ\tilde{\Delta}(\Lambda(y))\right\rangle.$
Using the formula (4.2), we will now show that when $\mathbb{G}$} is of Kac type and $a\in\operatorname{Pol}({\mathbb{G}})$ then $E(a)\in L^{\infty}({\mathbb{G}})$. From this it follows that $E$ preserves $L^{\infty}({\mathbb{G}})$, since $E$ is normal and $\operatorname{Pol}({\mathbb{G}})$ is ultraweakly dense in $L^{\infty}({\mathbb{G}})$, and hence that $E$ maps $L^{\infty}({\mathbb{G}})$ to $L^{\infty}_{z}({\mathbb{G}})$ by Lemma 4.1. Since $h$ is a trace, the map $\Lambda(x)\rightarrow\Lambda(x^{*})$ extends to an anti-unitary $J$ on $L^{2}({\mathbb{G}})$ with the property that $JL^{\infty}({\mathbb{G}})J=L^{\infty}({\mathbb{G}})^{\prime}$. By von Neumann’s bicommutant theorem and density of $\operatorname{Pol}({\mathbb{G}})$ in $L^{\infty}({\mathbb{G}})$, it therefore suffices to show that $E(a)$ commutes with $JbJ$ for all $b\in\operatorname{Pol}({\mathbb{G}})$. To see this, fix again $x,y\in\operatorname{Pol}({\mathbb{G}})$ and note that $JbJ\Lambda(x)=\Lambda(xb^{*})$. Using (4.2), we now compute as follows:
@@
$=\left\langle\Lambda\otimes\Lambda(a_{(2)}x_{(1)}\otimes a_{(1)}x_{(2)}),\Lambda\otimes\Lambda(y_{(1)}b_{(1)}\otimes y_{(2)}b_{(2)})\right\rangle$